OpenStudy (anonymous):

-Under the Quadratic equation (ROOTS) Determine the possible values of k so that the roots of 4x^2kx+35=0 differ by 1

OpenStudy (anonymous):

Is there meant be be a plus or a minus before the kx?

OpenStudy (anonymous):

yes that was what i was going to ask too!!

OpenStudy (anonymous):

If there isn't, I'm way out of my league :P

OpenStudy (anonymous):

Anyway, you'll end up with something along the lines of \[x=\frac{-k \pm \sqrt{k^2-c}}{8}\], and you need to say that the difference between the two roots (two because of the \(\pm\) sign) is 1. That should let you find k.

OpenStudy (anonymous):

\[4x^2kx+35=0 \]

OpenStudy (anonymous):

that's the equation

OpenStudy (anonymous):

I see no way of solving it in that case!

OpenStudy (anonymous):

I feel there is a misprint in yr book. it shud be + kx + 35

OpenStudy (anonymous):

ur ryt.. \

OpenStudy (anonymous):

there is a plus sign b4 the kx

OpenStudy (anonymous):

let me check now......

OpenStudy (anonymous):

sorry, can't seem to get it right....☺

OpenStudy (anonymous):

let the two roots be alpha and beta alhpha + beta = -k/4 (alpha)(beta)=35/4 (alpha + beta ) ^2 = ( alpha - beta )^2 + 4(alpha)(beta) k^2/16 = 1+35 k^2 = 36(16) k = 6* 4 = 24

OpenStudy (anonymous):

is my anwer correct ?

OpenStudy (anonymous):

yes Abhinav yr ans is correct !! Congrats n a medal ...☺